### v1

# The chromatic number of finite projective spaces

Abstract
The chromatic number of the finite projective space $\mathrm{PG}(n-1,q)$, denoted $\chi_{q}(n)$, is the minimum number of colors needed to color its points so that no line is monochromatic. We establish a recursive upper bound $\chi_{q}(n)\leq\chi_{q}(d)+\chi_{q}(n-d)$ for all $1\leq d<n$ and use it to prove new upper bounds on $\chi_{q}(n)$ for all $q$.
For $q=2$, we further refine this recursion and prove that
$\chi_{2}(n)\leq\lfloor 2n/3\rfloor+1$
for all $n\geq 2$, and that this bound is tight for all $n\leq 7$. In particular, this recovers all previously known cases for $n\leq 6$ and resolves the first open case $n=7$. On the lower-bound side, using a connection with multicolor Ramsey numbers for triangles, we note that
$\chi_{2}(n)\geq(1-o(1))\,\frac{n}{\log n}.$
We also consider $\chi_{q}(t;n)$, the minimum number of colors needed to color the points of $\mathrm{PG}(n-1,q)$ with no monochromatic (t1)(t-1)-dimensional subspace, and establish an equivalence between $\chi_{q}(t;n)$ and the multicolor vector-space Ramsey numbers $R_{q}(t;k)$. Using this equivalence together with our upper bounds on $\chi_{q}(t;n)$, we improve, for every fixed $t$, the best known lower bounds on $R_{q}(t;k)$ from $\Omega_{q}(\log k)$ to $\Omega(k)$. Finally, we improve the upper bound on $R_{2}(2;k)$ from a tower function of height $k$ to the function $k\log k$.
1 Introduction
Let $\mathbb{F}_{2}^{n}$ denote the $n$-dimensional vector space over $\mathbb{F}_{2}$ and let $\mathrm{PG}(n-1,2)$ denote the corresponding (n1)(n-1)-dimensional projective space. We identify the points of $\mathrm{PG}(n-1,2)$ with the non-zero vectors of $\mathbb{F}_{2}^{n}$. A line in $\mathrm{PG}(n-1,2)$ then corresponds to a triple $\{x,y,x+y\}$ with distinct $x,y\in\mathbb{F}_{2}^{n}\setminus\{0\}$. These lines define a $3$-uniform hypergraph $(V_{n},E_{n})$ with,
$V_{n}=\mathbb{F}_{2}^{n}\setminus\{0\}\quad\text{and}\quad E_{n}=\bigl\{\{x,y,x+y\}:x,y\in V_{n},\ x\neq y\bigr\},$
whose chromatic number is called the chromatic number $\chi_{2}(n)$ of the binary projective space. This is the least number of colors needed to color the points of PG(n1,2)PG(n-1,2) such that no line is monochromatic. Note that a color class contains no three collinear points; that is, each color class is a (projective) cap (see [16] for a recent survey on caps). Therefore, $\chi_{2}(n)$ is also the minimum number of caps that partition $\mathrm{PG}(n-1,2)$. Equivalently, $\chi_{2}(n)$ is the smallest number of sum-free sets in the abelian group $\mathbb{F}_{2}^{n}$ needed to partition the set of non-zero elements.
It can be easily shown that $\chi_{2}(2)=2$. It is a standard exercise in the context of Property B and minimal non-2-colorable hypergraphs (see, for example, [31, Section 1.3]) that $\chi_{2}(3)=3$, that is, the chromatic number of the Fano plane is $3$. The next case is $\chi_{2}(4)=3$, which follows from the well-known partition of $\mathrm{PG}(3,2)$ into elliptic quadrics [9].
The chromatic number of binary projective spaces was studied by Rosa [22, 23] in his investigation of more general Steiner triple systems (see [8, 5] for further results in this direction). He showed $\chi_{2}(5)=4$, and asked if $\chi_{2}(6)=5$; this was confirmed by Fugère, Haddad and Wehlau [12]. As a corollary of our Theorem 1, we obtain $\chi_{2}(7)=5$, thus solving the next open case.
In this paper, we prove new bounds on $\chi_{2}(n)$, recovering all the small cases proved earlier and improving the previous general upper bound. By coloring the complement of a hyperplane with one color and recursively iterating inside the hyperplane, it can be easily shown that $\chi_{2}(n)\leq n$. For example, let $i=1,\dots,n$, define
$C_{i}=\{x=(x_{1},\dots,x_{n})\in\mathbb{F}_{2}^{n}\setminus\{0\}:x_{j}=0\ \text{for all }j<i,\ \text{and }x_{i}=1\}.$
Then $PG(n-1,2)=\bigsqcup_{i=1}^{n}C_{i}$, and each $C_{i}$ is a cap. This bound can be improved to
$\chi_{2}(n)\leq n-1\ \text{for all}\ n\geq 4,$
using the recursion $\chi_{2}(n)\leq\chi_{2}(n-1)+1$ that follows from the above argument, together with the well-known base case of $n=4$.
Since $\chi_{2}(4)$, $\chi_{2}(5)$, and $\chi_{2}(6)$ are $3$, 44, and $5$, respectively, Haddad conjectured in [15] that $\chi_{2}(n)$ must increase by $1$ when the dimension increases by $1$ for all $n\geq 3$. Rosa [24] mentions that this conjecture was disproved by Blokhuis. We were unable to find the argument of Blokhuis, but the recursion $\chi_{2}(n)\leq\chi_{2}(d)+\chi_{2}(n-d)$ that follows from taking quotient spaces (we prove this in a more general setting in Lemma 1), combined with $\chi_{2}(4)=3$, implies that $\chi_{2}(n)\leq 3n/4+O(1)$ (see Corollary 1). We prove the following improvement to this upper bound.
Theorem 1.
For all $n\geq 2$
$\chi_{2}(n)\leq\lfloor 2n/3\rfloor+1.$
We will see below that this upper bound is tight for all $n\leq 7$. The key ingredient in the proof of Theorem 1 is the following improved recursion. For $1\leq d\leq n$,
$\chi_{2}(n)\leq\chi_{2}(d)+\chi_{2}(n+1-d)-1.$ (1)
Starting from the base case of $\chi_{2}(4)=3$, this yields Theorem 1. See Section 3 for the proof of the recursion, which uses a “reserved color” and a one-bit “lift”. This mechanism is in the same spirit as the “templates”-based approach for Schur numbers [25, 2].
For lower bounds on $\chi_{2}(n)$, we use a connection with multicolor Ramsey numbers for triangles. Recall that $R(3;k)$ denotes the smallest $N$ such that any $k$-coloring of the edges of a complete graph $K_{N}$ yields a monochromatic triangle. Around a decade before Ramsey’s foundational work, Schur [27] implicitly showed that $\Omega(2^{k})\leq R(3;k)\leq O(k!)$. Erdős later conjectured that $R(3;k)=2^{\Theta(k)}$. We refer to [18] for a historical account.
If we could properly color $\mathrm{PG}(n-1,2)$ with $k$ colors, then by identifying the $2^{n}$ vectors of $\mathbb{F}_{2}^{n}$ with the vertices of a $K_{2^{n}}$, we could color each edge $\{u,v\}$ of $K_{2^{n}}$ by the color of the difference $(v-u)\in\mathbb{F}_{2}^{n}\setminus\{0\}$, identified as a point of $\mathrm{PG}(n-1,2)$. So, a proper $k$-coloring of $\mathrm{PG}(n-1,2)$ would produce a $k$-edge-coloring of $K_{2^{n}}$ with no monochromatic triangles, and thus
$\chi_{2}(n)\leq k\implies R(3;k)>2^{n}.$ (2)
Our upper bound in Theorem 1 then implies that
$R(3;k)>2^{3(k-1)/2}>2.828^{k-1}.$ (3)
This connection between partitioning an abelian group into sum-free sets and $R(3;k)$ goes back to Abbott and Hanson [1] and ultimately Schur [27].
In the other direction, using eq. 2 and the classical asymptotic upper bound [14] of $R(3;k)\leq ek!+1$, we get the following lower bound on $\chi_{2}(n)$. All logarithms in this paper have base $2$.
Proposition 1.
$\chi_{2}(n)\geq(1-o(1))\,\frac{n}{\log n}.$
This highlights a wide gap: current lower bounds for $R(3;k)$ are exponential in $k$ [25, 2], while the best upper bounds are super-exponential. Any sublinear improvement for $\chi_{2}(n)$ would push these lower bounds of $R(3;k)$ beyond the exponential regime, which would disprove the conjecture of Erdős. At a concrete level, the only exact values known are $R(3;1)=3$, $R(3;2)=6$, and $R(3;3)=17$. For $k=4$, the current status is $51\ \leq\ R(3;4)\ \leq\ 62$ with the lower bound due to Chung [7], the upper bound due to Fettes, Kramer and Radziszowski [10], improving the earlier $R(3;4)\leq 64$ of Sánchez-Flores [26]; it is conjectured in [30] that $R(3;4)=51$ (see also the dynamic survey [19]). We will see below that these small cases can be used to find good bounds on $\chi_{2}(n)$.
Since the seminal upper bound of Greenwood–Gleason [14], progress in the general upper bounds of $R(3;k)$ has only focused on improving the leading constant in front of $k!$. The connection to $R(3;k)$ highlights that the chromatic number of (binary) projective spaces is a difficult quantity to pin down, exactly or even asymptotically. Nevertheless, we hope that the new upper bound of Theorem 1 opens the door to further improvements in the asymptotics of $R(3;k)$.
Beyond the binary case, we also study chromatic number of projective spaces over general finite fields and allow higher-dimensional forbidden subspaces. For a prime power $q$ and an integer $t\geq 2$ we write $\chi_{q}(t;n)$ for the chromatic number of the hypergraph on $\mathrm{PG}(n-1,q)$ whose edges are the sets of points of (t1)(t-1)-dimensional projective subspaces. For $t=2$ this is the chromatic number $\chi_{q}(n)$ of $\mathrm{PG}(n-1,q)$ with respect to lines. In the general setting, we prove the following recursion in Lemma 1.
$\chi_{q}(t;n)\leq\chi_{q}(t;n-d)+\chi_{q}(t;d).$
Since $\chi_{q}(t;n)=1$ for all $n\leq t-1$, we obtain the following upper bound.
Theorem 2.
For all prime powers $q$ and integers $n,t\geq 2$,
$\chi_{q}(t;n)\leq\lceil n/(t-1)\rceil.$
For $t=2$, we use better base cases to obtain stronger bounds.
Theorem 3.
For all fixed prime powers $q\geq 5$,
$\chi_{q}(n)\ \leq\ n/2+O(1)$
and, for q=3and 4q=3\ $\text${and}\ 4,
$\chi_{q}(n)\ \leq\ 3n/5+O(1).$
We then show an equivalence between $\chi_{q}(t;n)$ and multicolor vector space Ramsey numbers $R_{q}(t;k)$. These Ramsey numbers were originally introduced in [13] (also see [28]) and have recently been studied in [11, 17]. Recall that $R_{q}(t;k)$ is the smallest integer $n$ such that in every $k$-coloring of the $1$-dimensional linear subspaces of $\mathbb{F}_{q}^{n}$, there is a monochromatic $t$-dimensional linear subspace. Equivalently, $R_{q}(t;k)$ is the smallest integer $n$ such that in every $k$-coloring of the points of $\mathrm{PG}(n-1,q)$ there exists a monochromatic (t1)(t-1)-dimensional projective subspace. Therefore,
$\chi_{q}(t;n)\leq k\iff R_{q}(t;k)>n.$
Using our results, we thus obtain new lower bounds for $R_{q}(t;k)$ and also significantly improve the upper bound on $R_{2}(2;k)$.
Theorem 4.
For every prime power $q\geq 2$ and all integers $t\geq 2,\ k\geq 1$
$R_{q}(t;k)>(t-1)k.$
Theorem 5.
For all integers $k\geq 1$
$R_{2}(2;k)=O(k\log k).$
This improves the lower bounds on $R_{q}(t;k)$ from $\Omega(\log k)$ to $\Omega(k)$, and the upper bound on $R_{2}(2;k)$ from a tower function of height $k$ (see [11]) to the function $k\log k$. Note that for $t=2$, the lower bound $R_{q}(2;k)>k$ follows from $\chi_{q}(n)\leq n$, and thus the improved upper bounds in Theorem 3 give better lower bounds on $R_{q}(2;k)$. In particular, $R_{q}(2;k)>2k-O(1)$ for all fixed $q\geq 5$.
The paper is organized as follows. In Section 2 we introduce the general chromatic numbers $\chi_{q}(t;n)$, establish the basic recursion, and obtain upper bounds on $\chi_{q}(n)$. In Section 3 we focus on the binary case $q=2$, prove the recursion in Equation 1, and derive Theorem 1. In Section 4 we develop the connection to classical multicolor Ramsey numbers for triangles and determine exact values of $\chi_{2}(n)$ for small $n$. In Section 5 we make the link to multicolor vector space Ramsey numbers $R_{q}(t;k)$ and use our bounds on $\chi_{q}(t;n)$ to improve both lower and upper bounds for $R_{q}(t;k)$. In Section 6 we discuss future work and related open problems.
2 General upper bounds on the chromatic number
We use the convention that all dimensions are projective. Recall that we obtain a projective space $\mathrm{PG}(n-1,\mathbb{F})$ from the $n$-dimensional vector space $V$ over the field $\mathbb{F}$ in the following way: The points, lines, $\ldots$, (d1)(d-1)-spaces of $PG(n-1,\mathbb{F})$ are the $1$-dimensional, 22-dimensional, $\dots$, $d$-dimensional vector subspaces of $V$. For $v=(v_{1},\dots,v_{n})\in\mathbb{F}_{q}^{n}\setminus\{0\}$, we write
$v]=[v_{1}:\dots:v_{n}]\in\mathrm{PG}(n-1,q)$
for the point corresponding to the $1$-dimensional subspace $\langle v\rangle\subset\mathbb{F}_{q}^{n}$. Thus, $v]=[\lambda v]$ for every $\lambda\in\mathbb{F}_{q}^{\times}$.
Fix a prime power $q$ and an integer $t\geq 2$. For each integer $n\geq 1$, let $\chi_{q}(t;n)$ denote the smallest $k$ such that the points of $\mathrm{PG}(n-1,q)$ can be colored with $k$ colors in such a way that no (t1)(t-1)-dimensional projective subspace is monochromatic. Equivalently, this is the chromatic number of the hypergraph whose vertices are the points of $\mathrm{PG}(n-1,q)$ and whose hyperedges are the sets of points of each (t1)(t-1)-dimensional projective subspace.
For $q=2$ and $t=2$, a color class is a cap (no three collinear), so a proper coloring is the same as a partition into caps. For $q\geq 3$ and $t=2$, a color class must avoid a full $(q+1)$-point line. In other words, its complement is a blocking set [3]. We denote $\chi_{q}(2;n)$ as $\chi_{q}(n)$. In particular, partitions into caps give upper bounds on $\chi_{q}(n)$, but the true value of $\chi_{q}(n)$ can be smaller.
We can prove the following upper bounds. Let $S$ be a projective subspace of $\mathrm{PG}(n-1,q)$ of dimension ntn-t. Then $S\cong\mathrm{PG}(n-t,q)$, so we can color $S$ properly with $\chi_{q}(t;n-t+1)$ colors. Color all points outside $S$ with one new color. Then no (t1)(t-1)-dimensional projective subspace is monochromatic. Hence, for all prime powers $q\geq 2$ and for $n\geq t\geq 2$,
$\chi_{q}(t;n)\ \leq\ \chi_{q}(t;n-t+1)+1.$ (4)
In fact, we can do better. We have the following recursion for $\chi_{q}(t;n)$ that gives an improved upper bound in general.
Lemma 1.
For all prime powers $q\geq 2$, for all integers $t\geq 2$, and for all integers $1\leq d\leq n-1$,
$\chi_{q}(t;n)\leq\chi_{q}(t;n-d)+\chi_{q}(t;d).$ (5)
Proof.
Represent the points of $\mathrm{PG}(n-1,q)$ as $v_{1}:\dots:v_{n}]$. Define
$A\;=\;\bigl\{[v_{1}:\dots:v_{n}]:v_{n-d+1}=\dots=v_{n}=0\bigr\},$
$B\;=\;\bigl\{[v_{1}:\dots:v_{n}]:v_{1}=\dots=v_{n-d}=0\bigr\}.$
Then $A\cong\mathrm{PG}(n-d-1,q)$ and $B\cong\mathrm{PG}(d-1,q)$. If $v=[v_{1}:\dots:v_{n}]\in\mathrm{PG}(n-1,q)$ is a point, then we write
$v=[a:b].$
The point $[a:b]$ lies in $A$ if and only if $b=0$.
Let $c_{A}$ be a proper coloring of $A$ with $\chi_{q}(t;n-d)$ colors, and let $c_{B}$ be a proper coloring of $B$ with $\chi_{q}(t;d)$ colors. Use disjoint color palettes for $c_{A}$ and $c_{B}$. Define a coloring $c$ of $\mathrm{PG}(n-1,q)$ by
$c([a:b])\;=\;\begin{cases}c_{A}([a:0]),&\text{if }b=0,\\[3.0pt] c_{B}([b]),&\text{if }b\neq 0.\end{cases}$
Thus, $c$ uses $\chi_{q}(t;n-d)+\chi_{q}(t;d)$ colors.
We show that no (t1)(t-1)-dimensional projective subspace is monochromatic. Let $S\subseteq\mathrm{PG}(n-1,q)$ be a (t1)(t-1)-dimensional projective subspace. Then $S$ must satisfy one of the three cases depending on the dimension of $S\cap A$.
  • Suppose that $S\cap A=\emptyset$. Then every point of $S$ has the form $[a:b]$ with $b\neq 0$, and hence $c([a:b])=c_{B}([b])$. Consider the projection $\varphi([a:b])=b$, which maps points of $\mathrm{PG}(n-1,q)$ to points of $\mathrm{PG}(d-1,q)$. We show that $\varphi\big|_{S}$ is injective, which would imply that $\varphi(S)$ is a monochromatic (t1)(t-1)-space in $B$; a contradiction.
    Say, $\varphi\big|_{S}$ is not injective, i.e., there exists $a_{1}:b_{1}]\neq[a_{2}:b_{2}]$ on $S$ such that $\varphi([a_{1}:b_{1}])=\varphi([a_{2}:b_{2}])$, and hence $b_{1}=\lambda b_{2}$, for some $\lambda\neq 0$. Then, the point with coordinates $a_{1}:b_{1}]-\lambda[a_{2}:b_{2}]=[a_{1}-\lambda a_{2}:0]$ lies in $S\cap A$, which is a contradiction.
  • If $S$ contains a point $x=[a_{1}:b_{1}]$ with $b_{1}=0$ and a point $y=[a_{2}:b_{2}]$ with $b_{2}\neq 0$, then $c(x)$ is a color from the palette of $c_{A}$ and $c(y)$ is a color from the palette of $c_{B}$. Since the palettes are disjoint, $c(x)\neq c(y)$.
  • If $S\subseteq A$, then the coloring $c$ on $S$ coincides with $c_{A}$. Since $c_{A}$ is proper for (t1)(t-1)-spaces in $A$, $S$ is not monochromatic.
In all cases $S$ is not monochromatic. Therefore, $c$ is a proper coloring of $\mathrm{PG}(n-1,q)$, and this gives
$\chi_{q}(t;n)\ \leq\ \chi_{q}(t;n-d)+\chi_{q}(t;d).$
Remark 1.
We can give a geometric coordinate-free explanation of the recursion as follows. Take an (nd1)(n-d-1)-dimensional projective subspace $A$ of PG(n1,q)PG(n-1,q), and let $B=PG(n-1,q)/A\cong PG(d-1,q)$ be the quotient space with respect to $A$. Color the points of $A$ according to a minimal proper coloring of $A$ and the points outside $A$ according to a minimal proper coloring of $B$, with a point $x$ outside $A$ in PG(n1,q)PG(n-1,q) being identified with the point $\langle x,A\rangle$ of $B$.
Corollary 1.
For all integers $n\geq 2$,
$\chi_{2}(n)\leq\lceil 3n/4\rceil.$
Proof.
Applying Lemma 1 with $q=t=2$, $d=4$ and $\chi_{2}(4)=3$, we have
$\chi_{2}(n)\leq 3+\chi_{2}(n-4)\qquad(n\geq 5).$
Write $n=4m+r$ with $r\in\{1,2,3,4\}$. Iterating the inequality $m$ times gives
$\chi_{2}(n)\leq 3m+\chi_{2}(r).$
Using $\chi_{2}(r)=r$ for all $r\in\{1,2,3\}$ and $\chi_{2}(4)=3$, we get the stated bound. ∎
We now prove new bounds on $\chi_{q}(n)$ for $q>2$. A non-trivial blocking set [^1] A blocking set with respect to projective lines in $\mathrm{PG}(n-1,q)$ is a set of points that has nonempty intersection with every line of $\mathrm{PG}(n-1,q)$. A blocking set that contains a line is called trivial. exists for $\mathrm{PG}(2,q)$, $q>2$ [21]. Coloring the points of this blocking set with one color and the remaining points with another color gives a proper $2$-coloring, showing that $\chi_{q}(3)\leq 2$ for all $q>2$. Since $\chi_{q}(2)=2$ and $\chi_{q}(3)\geq\chi_{q}(2)$, we have $\chi_{q}(3)=2$ for all $q>2$. Applying Lemma 1 with $d=3$ and iterating shows $\chi_{q}(n)\ \leq\ 2n/3+O(1)$.
In fact, we can do better.
Theorem 3.
For all fixed prime powers $q\geq 5$,
$\chi_{q}(n)\ \leq\ n/2+O(1)$
and, for q=3and 4q=3\ $\text${and}\ 4,
$\chi_{q}(n)\ \leq\ 3n/5+O(1).$
Proof.
For $q\geq 5$, non-trivial blocking sets in $\mathrm{PG}(3,q)$ exist (see [29] for $q\geq 7$ and [20] for $q=5$). As in the case of $\mathrm{PG}(2,q)$ above, any non-trivial blocking set yields a proper $2$-coloring, so $\chi_{q}(4)\leq 2$. Since $\chi_{q}(4)\geq 2$, we have $\chi_{q}(4)=2$. Applying Lemma 1 with $d=4$ and iterating gives the first result.
For $q=3$ and $q=4$, non-trivial blocking sets in $\mathrm{PG}(3,q)$ do not exist (for the proof of non-existence see [29] for $q=3$ and [6] for $q=4$). If there were a $2$-coloring of $\mathrm{PG}(3,q)$, then it would give a non-trivial blocking set, so $\chi_{3}(4)>2$ and $\chi_{4}(4)>2$. Our computations [4] produced $3$-colorings of $\mathrm{PG}(4,3)$ and $\mathrm{PG}(4,4)$. Thus, $\chi_{3}(5)\leq 3$ and $\chi_{4}(5)\leq 3$. Since $\chi_{q}(4)\leq\chi_{q}(5)$, we obtain $\chi_{3}(4)=\chi_{3}(5)=\chi_{4}(4)=\chi_{4}(5)=3$. Applying Lemma 1 with $d=5$ and iterating then gives the second result. ∎
3 Improved upper bounds for binary projective spaces
As we saw in Corollary 1, $\chi_{2}(n)\leq 3n/4+O(1)$. In this section, we improve this upper bound by proving the following recursion in this setting.
Lemma 2.
For every integer $1\leq d\leq n-1$,
$\chi_{2}(n)\leq\chi_{2}(d)+\chi_{2}(n+1-d)-1.$
[…]
Figure 1: Coordinate slice of $\mathrm{PG}(n-1,2)$: AA (first ndn\!-\!d coordinates) and $U$ (last $d$).
[…]
Figure 2: Reserved color $r$ in $U$ (vertical strip) and routing to $A^{+}$ via the extra coordinate $t$.
Proof.
Let $n-1\geq d\geq 1$. Represent the points of $\mathrm{PG}(n-1,2)$ as $v_{1}:\dots:v_{n}]$. Define
$A\;=\;\bigl\{[v_{1}:\dots:v_{n}]:v_{n-d+1}=\dots=v_{n}=0\bigr\},$
$U\;=\;\bigl\{[v_{1}:\dots:v_{n}]:v_{1}=\dots=v_{n-d}=0\bigr\}.$
Then $A\cong\mathrm{PG}(n-d-1,2)$ and $U\cong\mathrm{PG}(d-1,2)$. If $v=[v_{1}:\dots:v_{n}]\in\mathrm{PG}(n-1,2)$ is a point, then we write $v=[a:u]$ (see Figure 1).
Take a proper $\chi_{2}(d)$-coloring $c_{U}$ of $U\cong\mathrm{PG}(d-1,2)$ and fix a reserved color $r$ of $c_{U}$. Define a function
$t:\mathrm{PG}(n-1,2)\longrightarrow\{0,1\}$
by
$t([a:u])\;=\;\begin{cases}0,&u=0,\\ 1,&u\neq 0\text{ and }c_{U}([u])=r,\\ 0,&u\neq 0\text{ and }c_{U}([u])\neq r.\end{cases}$
So $t([a:u])$ depends only on $u]\in U$. [^2] The value of $t$ in the third case will not be used later; we set it to 0 only to make $t$ a well defined function on $\mathrm{PG}(n-1,2)$. Now extend $A$ to $A^{+}\cong\mathrm{PG}(n-d,2)$ by adding an extra coordinate (see Figure 2). Take a proper $\chi_{2}(n+1-d)$-coloring $c_{A^{+}}$ of $A^{+}$, using a palette disjoint from the colors of $c_{U}$. Define
$c([a:u])\;=\;\begin{cases}c_{U}([u]),&\text{if }u\neq 0\text{ and }c_{U}([u])\neq r,\\[3.0pt] c_{A^{+}}\bigl([a:t([a:u])]\bigr),&\text{otherwise}\end{cases}$
Thus, $r$ is never used, and the total number of colors is $\chi_{2}(d)-1+\chi_{2}(n+1-d)$.
We show that $c$ is proper. Assume a line $\ell\subset\mathrm{PG}(n-1,2)$ is monochromatic in color $\alpha$. As the palettes of $c_{U}$ and $c_{A^{+}}$ are disjoint, there are two cases.
Case 1: $\alpha$ is a color of $c_{U}$. Then by the definition of $c$, for every $x=[a:u]\in\ell$ we have $u\neq 0$ and $c_{U}([u])\neq r$. Consider the projection $\varphi([a:u])=u$. Say $\varphi\big|_{\ell}$ is not injective, i.e., there exists $a_{1}:u_{1}]\neq[a_{2}:u_{2}]$ on $\ell$ such that $\varphi([a_{1}:u_{1}])=\varphi([a_{2}:u_{2}])=u$, and hence $u_{1}=u_{2}=u$. The third point on $\ell$ is then $a_{1}:u_{1}]+[a_{2}:u_{2}]=[a_{1}+a_{2}:0]$. This point must be colored using the palette of $c_{A^{+}}$, which is a contradiction. Therefore, projecting $\ell$ to $U$ via $a:u]\mapsto[u]$ yields a line of $U$ whose three points are all colored $\alpha$ by $c_{U}$. This is again a contradiction.
Case 2: $\alpha$ is a color of $c_{A^{+}}$. Then for every $x=[a:u]\in\ell$ we have $u=0$ or $c_{U}([u])=r$ and $c(x)=c_{A^{+}}([a:t(x)])=\alpha$. Let $x_{1}=[a_{1}:u_{1}]$, $x_{2}=[a_{2}:u_{2}]$ and $x_{3}=[a_{1}{+}a_{2}:\,u_{1}{+}u_{2}]$ denote the three points on $\ell$. We will show that in all cases (i) $t(x_{3})=t(x_{1})+t(x_{2})$, and (ii) the points $y_{1}=[a_{1}:t(x_{1})],y_{2}=[a_{2}:t(x_{2})],$ and $y_{3}=[a_{1}{+}a_{2}:t(x_{3})]$ are distinct points in $A^{+}\cong\mathrm{PG}(n-d,2)$. This gives three collinear points $y_{1},y_{2}$, and $y_{3}$ in $A^{+}$, all receiving the same color $\alpha$ in the proper coloring $c_{A^{+}}$, a contradiction. Note that for $(ii)$ it will suffice to show that $y_{1}\neq y_{2}$.
The following are the three cases we need to consider.
  • Say both $u_{1}$ and $u_{2}$ are 0. Then $u_{1}{+}u_{2}=0$, and hence $t(x_{1})+t(x_{2})=0+0=0=t(x_{3})$. Since $x_{1}\neq x_{2}$, we must have $a_{1}\neq a_{2}$ and thus $y_{1}\neq y_{2}$.
  • Say exactly one of $u_{1},u_{2}$ is 0 and the other is $r$-colored. Then $u_{1}{+}u_{2}\neq 0$, and hence it must also be $r$-colored. Then, $t(x_{1})+t(x_{2})=1+0=1=t(x_{3})$. Since $t(x_{1})\neq t(x_{2})$, we have $y_{1}\neq y_{2}$.
  • Say both $u_{1}$ and $u_{2}$ are $r$-colored in the coloring $c_{U}$. If $u_{1}\neq u_{2}$, then we must have $u_{1}{+}u_{2}\neq 0$, and hence $u_{1},u_{2}$ and $u_{3}=u_{1}+u_{2}$ are three non-zero vectors with $c_{U}(u_{1})=c_{U}(u_{2})=c_{U}(u_{3})=r$, contradicting the fact that $c_{U}$ is proper. Therefore, $u_{1}=u_{2}$ and $u_{3}=u_{1}+u_{2}=0$. We again conclude that $t(x_{1})+t(x_{2})=1+1=0=t(x_{3})$. Moreover, $x_{1}\neq x_{2}$, but $u_{1}=u_{2}$, and hence we must have $a_{1}\neq a_{2}$, which again gives $y_{1}\neq y_{2}$.
Therefore, $c$ must be a proper coloring of $\mathrm{PG}(n-1,2)$, and thus
$\chi_{2}(n)\;\leq\;(\chi_{2}(d)-1)+\chi_{2}(n+1-d)\;=\;\chi_{2}(d)+\chi_{2}(n+1-d)-1.\qed$
Theorem 1.
For all $n\geq 2$
$\chi_{2}(n)\leq\lfloor 2n/3\rfloor+1.$
Proof.
Applying Lemma 2 with $d=4$ and $\chi_{2}(4)=3$, we have
$\chi_{2}(n)\leq 2+\chi_{2}(n-3)\qquad(n\geq 5).$
Write $n=3m+r$ with $r\in\{1,2,3\}$. Iterating the inequality $m$ times gives
$\chi_{2}(n)\leq 2m+\chi_{2}(r).$
Using $\chi_{2}(r)=r$ for all $r\in\{1,2,3\}$, we get the stated bound. ∎
4 Connection to classical Ramsey theory
Here we prove the lower bound on $\chi_{2}(n)$ given in Proposition 1. The same proof works for any finite abelian group $G$ after fixing a strict linear order on $G$: for each unordered pair $\{u,v\}$ with $u<v$, color the edge by the color of vuv-u. If the vertex-coloring of $G$ has no monochromatic solution to $x+y=z$ (i.e., each color class is sum-free), then the edge-coloring of the complete graph $K_{|G|}$ has no monochromatic triangles. This argument goes back to Abbott and Hanson [1] who generalized Schur’s idea [27]. For completeness, we make the connection with $\chi_{2}(n)$ precise and provide a detailed proof. All logarithms in the statement and the proof have base two.
Proposition 1.
$\chi_{2}(n)\geq(1-o(1))\bigl(n/\log n\bigr)$.
Proof.
We work in $\mathbb{F}_{2}^{n}$. Assume $\chi_{2}(n)\leq k$. Then there is a coloring $c:\mathbb{F}_{2}^{n}\setminus\{0\}\to[k]$ such that no triple $\{x,y,x+y\}$ is monochromatic. (Over $\mathbb{F}_{2}$ these three elements are distinct, since $x=y$ would give $x+y=0$.)
Use $c$ to color the edges of $K_{2^{n}}$ on vertex set $\mathbb{F}_{2}^{n}$ as follows: for every edge $\{u,v\}$, color by c(vu)c(v-u). This is well-defined since $v-u\neq 0$. If $u,v,w$ formed a monochromatic triangle, then
c(vu)=c(wv)=c(wu),c(v-u)=c(w-v)=c(w-u),
but (vu)+(wv)=wu(v-u)+(w-v)=w-u in $\mathbb{F}_{2}^{n}$, which contradicts the assumption that $c$ has no monochromatic triple. Hence, this edge-coloring of $K_{2^{n}}$ has no monochromatic triangle, so
$R(3;k)\ >\ 2^{n}.$ (6)
We take the contrapositive. If $R(3;k)\leq 2^{n}$, then $\chi_{2}(n)>k$. Fix $\alpha\in(0,1)$ and set $k=\big\lfloor\alpha\,n/\log n\big\rfloor$. Recall that the classical upper bound of Schur and all subsequent improvements are of the form $R(3;k)=O(k!)$. Therefore
[…]
For all sufficiently large $n$, it follows that $R(3;k)\leq 2^{n}$ and hence $\chi_{2}(n)>k=\lfloor\alpha\,n/\log n\rfloor$. Since $\alpha\in(0,1)$ was arbitrary, $\chi_{2}(n)\geq(1-o(1))(n/\log n)$. ∎
4.1 Small values
We will use the known values and bounds [19] for the multicolor Ramsey numbers for triangles to obtain lower bounds on $\chi_{2}(n)$ for small $n$.
$R(3;1)=3,\qquad R(3;2)=6,\qquad R(3;3)=17,$
$51\leq R(3;4)\leq 62,\qquad 162\leq R(3;5)\leq 307\ ,\qquad 538\leq R(3;6)\leq 1838.$
These bounds give the following values of $\chi_{2}(n)$:
  • $n=5$: Since $R(3;3)=17\leq 2^{5}$, eq. 6 gives $\chi_{2}(5)>3$. By Theorem 1, $\chi_{2}(5)\leq 4$. Hence $\chi_{2}(5)=4$.
  • $n=6,7$: Since $R(3;4)\leq 62<2^{6}$, we get $\chi_{2}(7)\geq\chi_{2}(6)>4$. By Theorem 1, $\chi_{2}(6)=\chi_{2}(7)=5$
Corollary 2.
$\chi_{2}(7)=5$
Thus $n=7$ is the first case $n\geq 4$ where $\chi_{2}(n)\neq n-1$. These small values highlight the usefulness of the Ramsey connection and also show that we recover known results for $n=5$ [23] and $n=6$ [12].
We record the current small cases and the best bounds that we have obtained.
[…]
5 Connection to vector space Ramsey numbers
We now recall the multicolor vector space Ramsey numbers. Let $\mathcal{L}_{1}(V)$ denote the set of $1$-dimensional subspaces of a finite-dimensional vector space $V$ over $\mathbb{F}_{q}$. For integers $t_{1},\dots,t_{k}\geq 1$, the number $R_{q}(t_{1},\dots,t_{k})$ is defined to be the smallest $n$ such that for every coloring $f:\mathcal{L}_{1}(\mathbb{F}_{q}^{n})\to[k]$ there exist an index $i\in[k]$ and a $t_{i}$-dimensional subspace $U\leq\mathbb{F}_{q}^{n}$ such that all $1$-dimensional subspaces of $U$ receive color $i$.
In the case $t_{1}=\dots=t_{k}=t$ we write $R_{q}(t;k):=R_{q}(t,\dots,t)$. Thus, the existence of a monochromatic (t1)(t-1)-space in $\mathrm{PG}(n-1,q)$ is the same as the existence of a $t$-dimensional subspace $U\leq\mathbb{F}_{q}^{n}$ whose $1$-dimensional subspaces are all the same color. This gives an equivalence between $\chi_{q}(t;n)$ and the vector space Ramsey numbers $R_{q}(t;k)$. Let $q$ be a prime power, $t\geq 2$, $n\geq t$, and $k\geq 1$. Then
$\chi_{q}(t;n)\leq k\quad\Longleftrightarrow\quad R_{q}(t;k)>n.$ (7)
Equivalently,
$R_{q}(t;k)=\min\{n:\chi_{q}(t;n)>k\}.$
Frederickson and Yepremyan [11] note that standard probabilistic arguments give lower bounds for $R_{2}(t;k)$ of order $\Omega(2^{t}\log k/t)$, and the same argument over $\mathbb{F}_{q}$ shows that, for all $q$,
$R_{q}(t;k)=\Omega\!\left(\frac{q^{t}\log_{q}k}{t}\right).$
They further prove the following upper bounds for $R_{q}(t;k)$ when $q\in\{2,3\}$.
Theorem 6 ([11, Theorem 1.3]).
There exists a constant $C_{0}\approx 13.901$ such that for $\sigma_{2}=2$ and $\sigma_{3}=C_{0}$ the following holds. Let $q\in\{2,3\}$, let $k\geq 2$, and let $t_{k}\geq\cdots\geq t_{1}\geq 2$. Then $R_{q}(t_{1},\dots,t_{k})$ is at most a tower of height $\sum_{i=1}^{k-1}(t_{i}-1)+1$ of the form
$R_{q}(t_{1},\dots,t_{k})\leq\sigma_{q}^{\sigma_{q}^{\cdot^{\cdot^{\cdot^{\sigma_{q}^{3t_{k}}}}}}}.$
In particular, in the case $t_{1}=\dots=t_{k}=t$, the height of the tower is (k1)(t1)+1(k-1)(t-1)+1.
For general $t$, we can use Lemma 1 to get improved lower bounds for $R_{q}(t;k)$ for all $t\geq 2$ and fixed $q\geq 2$. Note that there is no (t1)(t-1)-dimensional projective subspace in $\mathrm{PG}(t-2,q)$, so every 1-coloring is proper. Hence $\chi_{q}(t;t-1)=1$.
Theorem 4.
For every prime power $q\geq 2$ and all integers $t\geq 2,k\geq 1$,
$R_{q}(t;k)>(t-1)k.$
In particular, $R_{q}(t;k)=\Omega_{t}(k)$.
Proof.
Applying Lemma 1 with $\chi_{q}(t;t-1)=1$ and iterating shows $\chi_{q}(t;n)\ \leq\ \lceil\tfrac{n}{t-1}\rceil$. For every integer $k\geq 1$, with equivalence (7) and choice n=(t1)kn=(t-1)k it follows that $R_{q}(t;k)>(t-1)k$. ∎
Remark 2.
Using Equation 7, the recursion in Lemma 1 can be equivalently phrased as
$R_{q}(t;k_{1}+k_{2})>R_{q}(t;k_{1})+R_{q}(t;k_{2}),$
which could be used directly to obtain these lower bounds.
For $t=2$, we provide a further improvements to the lower bound.
Lemma 3.
For every integer $k\geq 1$, the following holds.
  • If $q=2$, then
    $R_{2}(2;k)\ \geq\ \frac{3}{2}\,k-O(1).$
  • If $q=3$ or $q=4$, then
    $R_{q}(2;k)\ \geq\ \frac{5}{3}\,k-O(1).$
  • If $q\geq 5$, then
    $R_{q}(2;k)\ \geq\ 2k-O(1).$
Proof.
Assume that there exist constants $\alpha>0$ and $C$ such that
$\chi_{q}(2;n)\leq\alpha n+C\quad\text{for all }n\geq 2.$
Fix $k>C$. If
$n\leq\left\lfloor\frac{k-C}{\alpha}\right\rfloor,$
then $\alpha n+C\leq k$, so $\chi_{q}(2;n)\leq k$. By equivalence (7) with $t=2$ this implies $R_{q}(2;k)>n$. Hence
$R_{q}(2;k)\ \geq\ \frac{k}{\alpha}-O(1).$
For $q=2$ put $\alpha=2/3$ (Theorem 1), for $q=3$ and $q=4$ put $\alpha=3/5$, and for $q\geq 5$ put $\alpha=1/2$ (by Theorem 3). This gives the three stated bounds. ∎
We can also significantly improve the tower-type upper bound on this number for $q=t=2$.
Theorem 5.
$R_{2}(2;k)\leq k\log k+O(1)$.
Proof.
Let $n=R_{2}(2;k)-1$. Because there is a coloring of the points of PG(n1,2)PG(n-1,2) with $k$ colors and no monochromatic lines, (2) implies that
$2^{n}<R(3;k)\leq ek!+1\leq 2^{k\log k+\log e}.$
Since $n=R_{2}(2;k)-1$, this gives
$R_{2}(2;k)<\log R(3;k)+1<k\log k+\log e+1.$
6 Further directions
6.1 Binary projective spaces
The central problem is to determine the asymptotic behavior of $\chi_{2}(n)$.
$\text{Is }\ \chi_{2}(n)=\Theta(n)\ \text{ or }\ \chi_{2}(n)=o(n)\ ?$
Currently, we have
$\chi_{2}(n)\ =\ \Omega\Bigl(\frac{n}{\log n}\Bigr)\qquad\text{and}\qquad\chi_{2}(n)\ \leq\ \lfloor 2n/3\rfloor+1.$
Proving $\chi_{2}(n)=o(n)$ would be a major breakthrough. A sublinear growth of $\chi_{2}(n)$ would force much stronger lower bounds for multicolor triangle Ramsey numbers than is currently known. As a small step towards that goal, our recursion in Lemma 2 shows that a good bound $\chi_{2}(s)\leq t$ for small values of $s$ and $t$ implies stronger asymptotic bounds, as follows.
Proposition 2.
For all integers $n\geq d\geq 2$,
$\chi_{2}(n)\leq\left\lfloor\frac{\chi_{2}(d)-1}{d-1}\ n\right\rfloor+C_{d},$
where $1\leq C_{d}\leq\chi_{2}(d)-1$ is a constant depending only on $d$. Moreover, one can take $C_{4}=1$.
Corollary 3.
For every integer $d\geq 2$,
$\lim_{k\rightarrow\infty}R(3;k)^{1/k}\geq 2^{(d-1)/(\chi_{2}(d)-1)}.$
Proof.
Combine Proposition 2 with the implication in (2). ∎
Applying this to $d=4$ from Theorem 1, we obtain:
$\lim_{k\to\infty}R(3;k)^{1/k}\ \geq\ 2^{3/2}\ \approx\ 2.828.$
The best-known lower bound is
[…]
As a first concrete step, it is natural to improve the constant $2/3$ in the upper bound of $\chi_{2}(n)$.
$\chi_{2}(8)=5\ \Rightarrow\ \lim_{k\to\infty}R(3;k)^{1/k}\ \geq\ 2^{7/4}\ \approx\ 3.364,$
$\chi_{2}(13)\leq 8\ \Rightarrow\ \lim_{k\to\infty}R(3;k)^{1/k}\ \geq\ 2^{12/7}\ \approx\ 3.2813.$
Thus, proving $\chi_{2}(8)=5$ or $\chi_{2}(13)=8$ would improve the current best lower bound on the growth rate of $R(3;k)$.
Problem 1.
Determine whether $\chi_{2}(8)=5$ or $\chi_{2}(13)=8$.
6.2 Non-binary projective spaces
Problem 2.
Fix a prime power $q\geq 3$. Determine the asymptotic growth of $\chi_{q}(n)$. Is $\chi_{q}(n)=\Theta(n)$?
As a first step in this direction, one can try to improve the constants in our linear upper bounds. It will be interesting to improve the constant $\alpha_{q}$ such that
$\chi_{q}(n)\ \leq\ \alpha_{q}n+O(1).$
By Theorem 3, for $q\geq 5$ we have $\chi_{q}(n)\leq n/2+O(1)$ and for $q=3,4$ we have $\chi_{q}(n)\leq 3n/5+O(1)$.
Problem 3.
Develop $q$-ary analogs of the recursion $\chi_{2}(n)\leq\chi_{2}(d)+\chi_{2}(n+1-d)-1.$
If such recursions exist, then they may give better upper bounds on $\chi_{q}(n)$ for fixed $q$.
6.3 General setting of vector space Ramsey numbers
As we have shown, improving the bounds in the more general setting of $\chi_{q}(t;n)$, will have direct consequences on the vector space Ramsey numbers $R_{q}(t;k)$.
Problem 4.
Fix a prime power $q$ and an integer $t\geq 3$. Determine the asymptotic growth of $\chi_{q}(t;n)$. Is it true that $\chi_{q}(t;n)=\Theta_{t}(n)$?
If one improves the lower bounds on $\chi_{q}(t;n)$, then one also improves the upper bounds on vector space Ramsey numbers $R_{q}(t;k)$. The current upper bounds on $R_{q}(t;k)$ for $q=3$ have tower growth (see [11]), in contrast to linear lower bounds. Determining the exact value of $\chi_{q}(t;n)$ for small values of $q,t,$ and $n$, will also be interesting, as it can improve the general upper bounds via the recursion that we have established in our work.
Acknowledgements
The authors thank Ferdinand Ihringer and Sam Mattheus for valuable insights and discussions.

### v2

@@
 Using this equivalence together with our upper bounds on $\chi_{q}(t;n)$, we improve, for every fixed $t$, the best known lower bounds on $R_{q}(t;k)$ from $\Omega_{q}(\log k)$ to $\Omega(k)$.
-Finally, we improve the upper bound on $R_{2}(2;k)$ from a tower function of height $k$ to the function $k\log k$.
 1 Introduction
@@
 It can be easily shown that $\chi_{2}(2)=2$.
-It is a standard exercise in the context of Property B and minimal non-2-colorable hypergraphs (see, for example, [31, Section 1.3]) that $\chi_{2}(3)=3$, that is, the chromatic number of the Fano plane is $3$.
+It is a standard exercise in the context of Property B and minimal non-2-colorable hypergraphs (see, for example, [32, Section 1.3]) that $\chi_{2}(3)=3$, that is, the chromatic number of the Fano plane is $3$.
 The next case is $\chi_{2}(4)=3$, which follows from the well-known partition of $\mathrm{PG}(3,2)$ into elliptic quadrics [9].
@@
 $\chi_{q}(t;n)\leq k\iff R_{q}(t;k)>n.$
-Using our results, we thus obtain new lower bounds for $R_{q}(t;k)$ and also significantly improve the upper bound on $R_{2}(2;k)$.
+Using our results, we thus obtain new lower bounds for $R_{q}(t;k)$.
 Theorem 4.
@@
 $R_{q}(t;k)>(t-1)k.$
+This improves the lower bounds on $R_{q}(t;k)$ from $\Omega(\log k)$ to $\Omega(k)$.
+Note that for $t=2$, the lower bound $R_{q}(2;k)>k$ follows from $\chi_{q}(n)\leq n$, and thus the improved upper bounds in Theorem 3 give better lower bounds on $R_{q}(2;k)$.
+In particular, $R_{q}(2;k)>2k-O(1)$ for all fixed $q\geq 5$.
 Theorem 5.
@@
 $R_{2}(2;k)=O(k\log k).$
-This improves the lower bounds on $R_{q}(t;k)$ from $\Omega(\log k)$ to $\Omega(k)$, and the upper bound on $R_{2}(2;k)$ from a tower function of height $k$ (see [11]) to the function $k\log k$.
-Note that for $t=2$, the lower bound $R_{q}(2;k)>k$ follows from $\chi_{q}(n)\leq n$, and thus the improved upper bounds in Theorem 3 give better lower bounds on $R_{q}(2;k)$.
-In particular, $R_{q}(2;k)>2k-O(1)$ for all fixed $q\geq 5$.
+This improves the upper bound $R_{2}(2;k)=O(k\log^{7}k$) that follows from the more general bound given by Hunter and Pohoata [17, Theorem 1.4]), which was obtained using the breakthrough result of Kelley and Meka [18].
+[^1] From the discussion around [17, Theorem 4.1] one can also deduce the upper bound of $O(k\log k)$ on $R_{2}(2;k)$ but it is not stated explicitly.
 The paper is organized as follows.
@@
 In Section 4 we develop the connection to classical multicolor Ramsey numbers for triangles and determine exact values of $\chi_{2}(n)$ for small $n$.
-In Section 5 we make the link to multicolor vector space Ramsey numbers $R_{q}(t;k)$ and use our bounds on $\chi_{q}(t;n)$ to improve both lower and upper bounds for $R_{q}(t;k)$.
+In Section 5 we establish the link to multicolor vector space Ramsey numbers $R_{q}(t;k)$ and use our bounds on $\chi_{q}(t;n)$ to improve the lower bounds for $R_{q}(t;k)$.
 In Section 6 we discuss future work and related open problems.
@@
 $R_{q}(t;k)=\Omega\!\left(\frac{q^{t}\log_{q}k}{t}\right).$
-They further prove the following upper bounds for $R_{q}(t;k)$ when $q\in\{2,3\}$.
-Theorem 6 ([11, Theorem 1.3]).
-There exists a constant $C_{0}\approx 13.901$ such that for $\sigma_{2}=2$ and $\sigma_{3}=C_{0}$ the following holds.
-Let $q\in\{2,3\}$, let $k\geq 2$, and let $t_{k}\geq\cdots\geq t_{1}\geq 2$.
-Then $R_{q}(t_{1},\dots,t_{k})$ is at most a tower of height $\sum_{i=1}^{k-1}(t_{i}-1)+1$ of the form
-$R_{q}(t_{1},\dots,t_{k})\leq\sigma_{q}^{\sigma_{q}^{\cdot^{\cdot^{\cdot^{\sigma_{q}^{3t_{k}}}}}}}.$
-In particular, in the case $t_{1}=\dots=t_{k}=t$, the height of the tower is (k1)(t1)+1(k-1)(t-1)+1.
 For general $t$, we can use Lemma 1 to get improved lower bounds for $R_{q}(t;k)$ for all $t\geq 2$ and fixed $q\geq 2$.
@@
 For $q=2$ put $\alpha=2/3$ (Theorem 1), for $q=3$ and $q=4$ put $\alpha=3/5$, and for $q\geq 5$ put $\alpha=1/2$ (by Theorem 3). This gives the three stated bounds. ∎
-We can also significantly improve the tower-type upper bound on this number for $q=t=2$.
+The best upper bounds on $R_{q}(t;k)$ in general are tower functions of height (k1)(t1)+1(k-1)(t-1)+1 [11, Theorem 1.3].
+For $q=2$, Hunter and Pohoata [17] used the breakthrough result of Kelley and Meka [18] to show that $R_{2}(2;k)=O(k\log^{7}k)$.
+From the connection to $R(3;k)$ we immediately get the following.
 Theorem 5.
@@
 Acknowledgements
-The authors thank Ferdinand Ihringer and Sam Mattheus for valuable insights and discussions.
+The authors thank Ferdinand Ihringer, Sam Mattheus, and Liana Yepremyan for valuable insights and discussions.